Show that ¥(x,t)= Asin(wt+kx) describes a wave motion


1
Expert's answer
2021-02-16T12:14:42-0500

Answer:


Step 1

Given:

ψ(x,t)=Asin(wt+kx)


Step 2

Calculation:

ψ(x,t)=Asin(wt+kx)

ψx,t=Asinwt+kx


We know the wave equation is


"∂^2ψ \over ∂x^2" = "1 \over v^2" "∂^2ψ \over ∂t^2" →(1)


let us find out all the derivatives 


"∂ψ \over ∂x" =Ak cos(wt+kx)


Again differentiate with respect to x,


"∂^2ψ \over ∂x^2" = Ak2  sin(wt+kx)(Sinceψ(x,t)=Asin(wt+kx)


"∂^2ψ \over ∂x^2" =−k2ψ(x,t)→(2)


Step 3

Now,Let's find out time derivatives 


"∂ψ \over ∂t" =Aw cos(wt+kx)


Again differentiate with respect to x,



"∂^2ψ \over ∂t^2" =−Aw2 sin(wt+kx) (Sinceψ(x,t)=Asin(wt+kx)



"∂^2ψ \over ∂t^2" = −wψ(x,t)→(3)


Put the values in(1) equation (2) & (3)


−k2 ψ(x,t)= "1 \over v^2" −w2 ψ(x,t)


Step 4

v2 ψ(x,t)= "w^2 \over k^2" ψ(x,t)

(Since one dimensional wave equationv= "w \over k" and v2 = "w^2 \over k^2" )



v2 ψ(x,t)=v2 ψ(x,t)


Hence it is proved.